Evolution and Scientists. A professor claims that more than 81% of all scientists believe in non-theistic evolution. In a 2014 Pew Research survey of a representative sample of 190 scientists connected to the American Association for the Advancement of Science (AAAS), 90% of these scientists (171 scientists) say they believe in non-theistic evolution. Test the professors claim using these data (a) Select the correct null and alternative hypotheses. Ho: p = 0.9 Hj: p > 0.9 0H0: p = 0.81 Hj: p > 0.81 Hyp 0.81 Ho: p = 0.9 Or?o: p-0.81 H: p<0.81 H: p <0.9
Answer #1

a) option B Hypothesis is Correct

b) z=(\frac{0.90-0.81}{\sqrt{\frac{0.81*0.19}{190}}})

z = 3.16

P value = P(z>3.16)

=0.000

c) Option B is Correct